OpenStudy (mtalhahassan2):

Determine the equation of the axis of symmetry of the parabola with points (-5,3) and (3,3) equally distant from the vertex on either side of it.

OpenStudy (aum):

Find the x-coordinate of the midpoint. The axis will be a vertical line passing through the midpoint.

OpenStudy (mtalhahassan2):

y= a(x-r) (x-s)

OpenStudy (mtalhahassan2):

So we have to used this formula

OpenStudy (aum):

What is the midpoint of (-5,3) and (3,3) ?

OpenStudy (mtalhahassan2):

1

OpenStudy (mtalhahassan2):

5-3 divide by 2

OpenStudy (aum):

midpoint = ( (x1+x2)/2, (y1+y2)/2 )

OpenStudy (mtalhahassan2):

what is that?

OpenStudy (mtalhahassan2):

( (x1+x2)/2, (y1+y2)/2 )

OpenStudy (aum):

Midpoint = \(\Large\left(\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2} \right) \) x-coordinate of midpoint = (-5+3)/2 = -2/2 = -1 x = -1 is the axis of symmtery.

OpenStudy (mtalhahassan2):

oh ok

OpenStudy (mtalhahassan2):

thanks buddy

OpenStudy (aum):

yw.